# Enhancing User Interest based on Stream Clustering and Memory Networks in Large-Scale Recommender Systems

### v5

@@ Abstract.
We conduct extensive offline and online experiments in a large-scale industrial RS. The results demonstrate that our model substantially outperforms other existing approaches, especially for users with sparse interest. UIE has been deployed in several large-scale RSs at Tencent since 2022, which was made public on 21 May 2024. In addition, UIE-based methods have also been successfully applied in candidate generation, pre-ranking, and context-DNN stages. Multiple teams have developed solutions based on UIE, focusing primarily on updating clustering algorithms and attention mechanisms. As far as we know, UIE has been deployed by many companies. The thoughts of UIE, dynamic streaming clustering and similarity enhancement, have inspired subsequent relevant works.
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; May 21, 2024; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’24), May 21, 2024, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/24/05doi: 10.1145/3690624.3709429ccs: Information systems Recommender systemsccs: Information systems Personalization
@@ 3.2. User Profile Enhancement (UPE)
Finally, generating user profile enhancement vector based on the most similar centroids and personal vector. After obtaining the ${K_{1}}$ most similar vectors, the weighted vector $\upsilon_{weighted}$ is calculated using attention mechanism, as shown in Formula 2. Furthermore, to achieve better personalized enhancement, a short length vector for each user is predefined called personal vector, which is used to achieve personalized fine-tuning of weighted vector. The weighted vector and personal vector are concatenated and used as the input of the generalization layer, as shown in Figure 2.
@@ 6. Conclusion
In this work, we first highlight the problem that lots of users in large-scale RSs merely have sparse interest, resulting in poor recommendation performance for them. To solve this difficult problem, we propose a novel user interest enhancement solution called UIE for enhancing user profile and user history behavior sequences. The key idea of UIE is to construct enhancement vectors for enhancing user interest from different perspectives, using the most similar centroids and the most relevant centroids which are generated by stream clustering and stored in different memory networks. By supplementing key information of user interest including user profile and user history behavior sequences, UIE remarkably improves the performance of ranking model on low-active users and achieves significant improvements on high-active users as well. For offline evaluation, we conduct extensive experiments to compare the performance of different models. The results show that UIE significantly exceeds other methods, especially on low-active users. Furthermore, we also conduct online A/B testing in a large-scale RS and the results demonstrate our model performs significantly better than other models, resulting in +7.96% UVC, +3.65% UDT on low-active users and +4.06% UVC, +2.19% UDT on high-active users. In addition, our model also improves diversity metrics on low-active and high-active users significantly. Similar methods are used for tail items as well to improve their exposure rate and recommendation performance. UIE has been deployed in multiple large-scale RSs and has inspired subsequent relevant works.

### v6

@@ Abstract.
We conduct extensive offline and online experiments in an industrial RS. The results demonstrate that UIE substantially outperforms other existing approaches, especially for users with sparse interest. UIE has been deployed in several large-scale RSs at Tencent since 2022, which was made public on 21 May 2024. In addition, UIE-based methods have also been successfully applied in candidate generation, pre-ranking, and context-DNN stages. Multiple teams have developed solutions based on UIE, focusing on updating clustering algorithms and attention mechanisms. As far as we know, UIE has been applied in multiple RSs, advertising systems and search engines. The thoughts of UIE, dynamic streaming clustering and similarity enhancement, have inspired subsequent relevant works.
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; May 21, 2024; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’24), May 21, 2024, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/24/05doi: 10.48550/arXiv.2405.13238ccs: Information systems Recommender systemsccs: Information systems Personalization
@@ 3.2. User Profile Enhancement (UPE)
Finally, generating user profile enhancement vector based on the most similar centroids and personal vector. After obtaining the ${K_{1}}$ most similar vectors, the weighted vector $\upsilon_{weighted}$ is calculated using attention mechanism, as shown in Formula 2. Furthermore, to achieve better personalized enhancement, a short length vector for each user is predefined called personal vector, which is used to achieve personalized fine-tuning of weighted vector. The weighted vector and personal vector are concatenated and used as the input of the generalization layer, as shown in Figure 2. We adopt several structures similar to PEPNet (ref3, ) as the generalization layer to further improve performance. The generalization layers of the following components also adopt the similar approach.
@@ 6. Conclusion
In this work, we first highlight the problem that lots of users in large-scale RSs merely have sparse interest, resulting in poor recommendation performance for them. To solve this difficult problem, we propose a novel user interest enhancement solution called UIE for enhancing user profile and user history behavior sequences. The key idea of UIE is to construct enhancement vectors for enhancing user interest from different perspectives, using the most similar centroids and the most relevant centroids which are generated by stream clustering and stored in different memory networks. By supplementing key information of user interest including user profile and user history behavior sequences, UIE remarkably improves the performance of ranking model on low-active users and achieves significant improvements on high-active users as well. For offline evaluation, we conduct extensive experiments to compare the performance of different models. The results show that UIE significantly exceeds other methods, especially on low-active users. Furthermore, we also conduct online A/B testing in a large-scale RS and the results demonstrate our model performs significantly better than other models, resulting in +7.96% UVC, +3.65% UDT on low-active users and +4.06% UVC, +2.19% UDT on high-active users. In addition, our model also improves diversity metrics on low-active and high-active users significantly. Similar methods are used for tail items as well to improve their exposure rate and recommendation performance. UIE has been deployed in multiple large-scale RSs and has inspired subsequent relevant works. Furthermore, the thoughts of UIE have also been applied in advertising systems and search engines.